November 19 2015
Attending: Team IU, Mat, Louise

- Follow up from DQ meeting: Joao reports that the cronjob to fill the isgoodrun metadata for FD files had indeed stopped. He restarted it an will work make this more robust.


Very good week of 13DB running.
Gap in data taking on Nov 17th appears to be opportunistic down time while beam was off.
On Monday 11/17 there were DAQ issues which led to multiple run starts in a matter of seconds and 2 hours of running with some fraction of "empty events". It looks like the issue was identified immediately and the lag was just in rousing an expert late at night.
Files size for the FD subruns was doubled on 11/13 and we see no ill effects from that. We do see fewer subruns per hour and more subrun to subrun consistantcy in our metrics as expected.

One oddity - the slow monopole trigger rate slowly increased between 11/15 and 11/16. No idea why. No other detector metrics show this. (Probably indicates a background of slow monopoles turned on recently ;) )


A new main nearline webpage is in development to handle the multiple trigger processing which is under study.
Justin and Ryan (from team IU) are also taking on webpage and plot making duties for the nearline effort.


Active number of FEBs stabilized after 11/16 DSO scan.
Maintenance expected to begin on Thursday.
Need update on ND good data keep up.
Long discussion on APD rates versus time... Fernanda is going to dust off her apd pixel rate tracking code.
Louise is going to setup a reference file for the ND onmon... working on an updating protocol.